OpenStudy (anonymous):

How many cups are there in 4.56 gallons? (Hint: 1 gallon = 4 quarts; 1 quart = 2 pints; 1 pint = 2 cups) A. 1.46 × 102 cups B. 7.30 × 101 cups C. 3.65 × 101 cups D. 1.82 × 101 cups E. 9.12 × 100 cups

OpenStudy (anonymous):

Can anyone please help?

OpenStudy (anonymous):

i think it's b but ....i got 73...i dont know what it means when it says times 101...

OpenStudy (anonymous):

r u sure that u have made no mistake in writing the question

OpenStudy (anonymous):

How many cups are there in 4.56 gallons? (Hint: 1 gallon = 4 quarts; 1 quart = 2 pints; 1 pint = 2 cups) A. 1.46 × 10^2 cups B. 7.30 × 10^1 cups C. 3.65 × 10^1 cups D. 1.82 × 10^1 cups E. 9.12 × 10^0 cups

OpenStudy (anonymous):

Sorry! It is like this! SORRY!

OpenStudy (anonymous):

:) now you can solve this... 1 gallon = 4 quarts = 4 * (2 pints) as 1 quart = 2 pint = 8 * (2 cups) as 1 pint = 2 cups this means 1 gallon = 16 cups. now, 4.56 gallons = ????????? i think you can calculate this one and you will get the answer what nuh said.

OpenStudy (whpalmer4):

4.56 gallons * 4 quarts / 1 gallon * 2 pints / 1 quart * 2 cups / 1 pint

OpenStudy (anonymous):

still finding it difficult ?? @RH

OpenStudy (anonymous):

If you ratio...so 4.56 times 4 times 2 times 2 and that equals 73 which is 7.3 times 10...so B

OpenStudy (anonymous):

so 4.56 x 4 x 2 ?

OpenStudy (anonymous):

@ANISH95

OpenStudy (anonymous):

4.56 * 4 * 2 * 2

OpenStudy (whpalmer4):

\[4.56 gal * \frac{4 qt}{1 gal}*\frac{2 pt}{1 qt}*\frac{2 c}{1 pt} = \]

OpenStudy (whpalmer4):

If you do it with the units present, you are protected against error — if you have it set up wrong, the units won't work out.

OpenStudy (anonymous):

@RH YOU MAY DO THE QUESTION THE WAY WHPALMER4 IS SAYING TOO... i hope you understood how to do it

OpenStudy (anonymous):

Yes I understand it now!! Thanks a lot! :)

OpenStudy (anonymous):

:) welcome